Solve by substitution:
2x+2y=2
3x-y=1

Answer by joecbaseball(37)


First, to do substitution, you need to solve for one of the variables in one of the equations. It does not matter which one. I will solve for y in the second equation.
Thus, 3x – y = 1 gives us y = 3x – 1.
Now, substitute this value for y into the first equation:
2x + 2y = 2
2x + 2(3x - 1) = 2
2x + 6x - 2 = 2
8x = 4
x = 4/8 = ½.
So, now that you have the value of x, you substitute that back into either equation to solve for y:
2x + 2y = 2
2(1/2) + 2y = 2
1 + 2y = 2. Solve for y.
2y = 1
y = ½.
So, your answers are x = ½, y = ½.
Good luck,
JoeC